Find the surface area of a sphere that has a volume of 288 pi ft cube. Give answer in terms of pi and as a decimal
Answer by sachi(548)

the volume of a sphere is 288 pi ft cube.
i.e. 4/3 pi r^3=288 pi
or r^3=288*3/4=216
or r=6
so the surface area of the cubu =4pi r^2=4 *pi*36=144pi ft sqr